# Scale-Invariant Gradient Aggregation for Constrained Multi-Objective Reinforcement Learning

### v1

@@ B.3.1 MORL to CMORL using the Lagrangian method
B.3.2 Quadratic Programming Problem
In the main text, we have proposed to obtain the aggregated gradient $\bar{g}_{\omega}^{\mathrm{ag}}$ by solving the quadratic programming (QP) problem (8), which can be addressed using an existing QP solver (Goldfarb & Idnani, 1983). However, due to the large dimensions of the policy parameter space where the problem is defined, it will take a long time to calculate the aggregated gradient. As an alternative, we propose to solve a Lagrange dual problem. Given that QP problems hold strong duality, the solution to the dual problem is equivalent to that of the original problem. The dual problem is then expressed as follows:
$\mathrm{max}_{\lambda\geq 0,\nu\geq 0}\;\mathrm{min}_{g}\;\frac{1}{2}g^{T}Hg+\sum_{i=1}^{N}\lambda_{i}(\omega_{i}e_{i}-g_{i}^{T}g)+\sum_{k\in I_{\mathrm{AC}}^{\epsilon}}\nu_{k}(b_{k}^{T}g+J_{C_{k}}(\theta_{\mathrm{old}})-d_{k}^{\prime})=:L(g,\lambda,\nu),$ (42)
where $\lambda$ and $\nu$ are Lagrange multipliers. Since $\mathrm{argmin}_{g}L(g,\lambda,\nu)$ is $H^{-1}(\sum_{i}\lambda_{i}g_{i}-\sum_{k}\nu_{k}b_{k})$, denoted by $g^{*}(\lambda,\nu)$, we can get the following problem by replacing $g$ with $g^{*}(\lambda,\nu)$:
$\mathrm{min}_{\lambda\geq 0,\nu\geq 0}\;$ $\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\lambda_{i}\lambda_{j}g_{i}^{T}H^{-1}g_{j}+\frac{1}{2}\sum_{i\in I_{\mathrm{AC}}^{\epsilon}}\sum_{j\in I_{\mathrm{AC}}^{\epsilon}}\nu_{i}\nu_{j}b_{i}^{T}H^{-1}b_{j}-\sum_{i=1}^{N}\sum_{j\in I_{\mathrm{AC}}^{\epsilon}}\lambda_{i}\nu_{j}g_{i}^{T}H^{-1}b_{j}$ (43)
$-\sum_{i=1}^{N}\lambda_{i}\omega_{i}e_{i}-\sum_{k\in I_{\mathrm{AC}}^{\epsilon}}\nu_{k}(J_{C_{k}}(\theta_{\mathrm{old}})-d_{k}^{\prime}),$
where $g_{i}^{T}H^{-1}g_{j}$, $b_{i}^{T}H^{-1}b_{j}$, and $g_{i}^{T}H^{-1}b_{j}$ can be computed in parallel using GPU. We have formulated the new QP problem (43) in the dual space of dimension $N+|I_{\mathrm{AC}}^{\epsilon}|\leq N+M$, which is much smaller than the dimension of the policy parameter space. Now, we solve (43) using a QP solver to get the optimal dual variables, which are denoted by $\lambda^{*}$ and $\nu^{*}$. Finally, the aggregated gradient can be expressed as follows using the optimal dual variables:
$\bar{g}_{\omega}^{\mathrm{ag}}=\sum_{i=1}^{N}\lambda^{*}_{i}H^{-1}g_{i}-\sum_{k\in I_{\mathrm{AC}}^{\epsilon}}\nu^{*}_{k}H^{-1}b_{k}.$ (44)
B.3.3 Network Architecture
The proposed method and baselines are based on the actor-critic framework, requiring policy and critic networks. Additionally, since the baselines use the Lagrangian method to handle the constraints, they use multiplier networks, as mentioned in Appendix B.3.1. We have implemented these networks as fully connected networks (FCNs), and their structures are presented in Table 5.
@@ B.3.3 Network Architecture
In the unconstrained tasks, MO-Gymnasium, we use standard critic networks that output scalar values for given observations and preferences. However, the standard critic networks usually estimate the objective or constraint functions with large biases, making it challenging to satisfy the constraints. Therefore, we use quantile distributional critics (Dabney et al., 2018) to lower the estimation biases in the Safety-Gymnasium and legged robot locomotion experiments. The quantile distribution critic outputs several quantile values of the discounted sum of the reward or cost functions for given observations and preferences. Then, the objective or constraint functions can be estimated by averaging the outputted quantiles. For the proposed method and LP3 that have similar frameworks to TRPO (Schulman et al., 2015), we use the TD($\lambda$) method (Kim et al., 2023) to train the distributional critic networks. For PD-MORL and CAPQL, the distributional critic networks are trained by reducing the Wasserstein distance between the current quantiles and the truncated one-step TD targets, as in TQC (Kuznetsov et al., 2020). The policy network, similar to other RL algorithms dealing with continuous action spaces (Schulman et al., 2015; Haarnoja et al., 2018; Fujimoto et al., 2018), outputs the mean and diagonal variance of a normal distribution. However, in the locomotion tasks, the policy networks are modified to output only the mean value by fixing the diagonal variance in order to lower the training difficulty.
B.3.4 Hyperparameter Settings
@@
$K$ 0
$G$infty
$H$ matrix - - Hessian of KL divergence Hessian of KL divergence
@@
Policy learning rate $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-5}$
@@
Local region size - - 0.14 0.1
$K$ 0
$G$infty
$H$ matrix - - Hessian of KL divergence Identity matrix
Appendix C Additional Experimental Results
In this section, we introduce and analyze the results of ablation studies on the hyperparameters of the proposed method. Next, the performance metrics of the experiment presented in Section 6.5, which compares the proposed method with the LS version, are reported. Lastly, we visualize the estimated CP fronts for tasks with two objectives.
C.1 Ablation Study on Hyperparameters
The proposed method has introduced new hyperparameters, $\epsilon$, $K$, and $G$. We first describe the role of each hyperparameter and show the results of the ablation study. $\epsilon$ is the local region size, where the local region is defined as $||\Delta\theta||_{H}\leq\epsilon$. Since $\epsilon$ adjusts the size of the policy update, it serves as the learning rate of the policy parameter. $K$ is introduced in (8) to mitigate the risk of constraint violation arising from linear approximation, and this is achieved by reducing the threshold by $K\epsilon$. Therefore, setting a large $K$ lowers the likelihood of constraint violations. However, assigning a too large value to $K$ can lead to an empty feasible region, $\{\theta|J_{C_{k}}(\theta)\leq d_{k}-K\epsilon\}=\emptyset$, causing the policy to break down. $G$ is introduced to scale down the policy gradient when $\alpha$ defined in (9) becomes large. $\alpha$ is defined as the ratio of the magnitude of the aggregated gradient $\bar{g}_{\omega}^{\mathrm{ag}}$ to that of each objective’s gradient. Consequently, $\alpha$ increases when the gradients of the objectives are in opposing directions, indicating a gradient conflict. Thus, $G$ serves to reduce the size of the policy gradient in such conflict situations. From this observation, a smaller $G$ implies a strict judgment of gradient conflicts, while a larger $G$ indicates a generous judgment of these conflicts.
Now, we introduce the ablation study on the hyperparameters. We conduct experiments on the quadrupedal task in the legged robot locomotion environment for various hyperparameter values. In particular, we also conduct the experiments at $K=0$, a setting that does not reduce the thresholds of the constraints, and $G=\infty$, a setting that does not consider gradient conflicts. The results of the ablation study are reported in Table 8, and trends in HV and SP values according to hyperparameter values are visualized in Figure 7. Table 8 confirms that the constraints are satisfied in all settings, and Figure 7 shows that the performance metric values do not change significantly across all settings. However, a significant performance drop is observed at $K=100$, which is likely because a huge $K$ causes the feasible region to become empty, causing the policy to be updated in the wrong direction, as mentioned earlier. We also see a slight decrease in HV for $G=1$, which is due to the strict judgment of gradient conflicts, making the policy updated conservatively. Except for these settings, there is no significant change in performance according to the hyperparameter settings, including $K=0$, which does not reduce the thresholds, and $G=\infty$, which does not consider gradient conflicts. From this observation, we recommend setting $K=0$ and $G=\infty$ to reduce the complexity of hyperparameter tuning. Also, we have conducted the main experiments with $K=0$ and $G=\infty$, as observed in Appendix B.3.4.
Table 8: Results of the ablation study on the hyperparameters. The reported values are obtained by averaging results from five random seeds.
Trust region size, $\epsilon$ $K$ $G$
Value $0.01$ $0.02$ $0.05$ $0.1$ $0.0$ $1.0$ $10.0$ $100.0$ $1$ $10$ $100$infty
HV (×$\times 10^{4}$) $\uparrow$ $4.736$ $4.676$ $4.840$ $4.765$ $4.839$ $4.915$ $4.920$ $0.000$ $4.064$ $4.829$ $4.872$ $4.839$
$\overline{\mathrm{SP}}$ (×$\times 10^{-2}$) $\downarrow$ $1.643$ $1.795$ $1.691$ $2.496$ $1.691$ $2.799$ $2.009$ $159.884$ $1.810$ $1.746$ $1.917$ $1.691$
Con1 ($\leq 0.025$) $0.005$ $0.005$ $0.015$ $0.012$ $0.015$ $0.006$ $0.008$ $0.024$ $0.003$ $0.004$ $0.008$ $0.015$
Con2 ($\leq 0.025$) $0.003$ $0.003$ $0.014$ $0.012$ $0.014$ $0.004$ $0.004$ $0.006$ $0.002$ $0.001$ $0.006$ $0.014$
Con3 ($\leq 0.4$) $0.322$ $0.326$ $0.338$ $0.340$ $0.338$ $0.324$ $0.320$ $0.329$ $0.300$ $0.291$ $0.325$ $0.338$
Figure 7: Trends in HV and SP according to hyperparameter settings.
C.2 Performance Metrics of Experiments in Section 6.5
We report the performance metric values of the experiments which compare the proposed method with the linear scalarization (LS) version, conducted in Section 6.5. The results are presented in Table 9.
Table 9: Experimental results comparing the proposed method with LS version. The reported values are obtained by averaging results from five random seeds.
CoMOGA (Ours) Linear Scalarization
Scale of the energy efficiency objective ×$\times 10.0$$\times 1.0$$\times 0.1$$\times 10.0$$\times 1.0$$\times 0.1$
HV (×$\times 10^{4}$) $\uparrow$ $5.823$ $5.941$ $5.886$ $5.205$ $5.942$ $4.690$
$\overline{\mathrm{SP}}$ (×$\times 10^{-2}$) $\downarrow$ $2.010$ $1.691$ $2.478$ $5.001$ $1.404$ $5.516$
Con1 ($\leq 0.025$) $0.016$ $0.015$ $0.012$ $0.006$ $0.007$ $0.017$
Con2 ($\leq 0.025$) $0.007$ $0.014$ $0.007$ $0.014$ $0.003$ $0.007$
Con3 ($\leq 0.4$) $0.336$ $0.338$ $0.326$ $0.340$ $0.309$ $0.315$
C.3 Visualization of Estimated CP Fronts
@@
As mentioned in Appendix B.2, we can obtain the estimated CP front from the trained universal policy. We use 20 equally spaced preferences to get the fronts for each task, and the obtained CP fronts are visualized in Figure 8. The hopper and Lunar-Lander tasks in the MO-Gymnasium are excluded from this visualization as they have three and four objectives, respectively.
(a) Quadruped
(b) Biped
@@
(e) Single-Point-Goal
(f) Single-Car-Goal
(g) Half-Cheetah
(h) Water-Reservoir
Figure 8: Visualization of Estimated CP Fronts. For each task, we plot the estimated fronts obtained from five random seeds in the same figure. The number of items in the fronts can be reduced due to constraint violations or non-dominant policies.

### v2

@@ C.3.1 MORL to CMORL using the Lagrangian method
C.3.2 Network Architecture
The proposed method and baselines are based on the actor-critic framework, requiring policy and critic networks. Additionally, since the baselines use the Lagrangian method to handle the constraints, they use multiplier networks, as mentioned in Appendix C.3.1. We have implemented these networks as fully connected networks (FCNs), and their structures are presented in Table 3.
@@ C.3.2 Network Architecture
In the unconstrained tasks, MO-Gymnasium, we use standard critic networks that output scalar values for given observations and preferences. However, the standard critic networks usually estimate the objective or constraint functions with large biases, making it challenging to satisfy the constraints. Therefore, we use quantile distributional critics [9] to lower the estimation biases in the Safety-Gymnasium and legged robot locomotion experiments. The quantile distribution critic outputs several quantile values of the discounted sum of the reward or cost functions for given observations and preferences. Then, the objective or constraint functions can be estimated by averaging the outputted quantiles. For the proposed method and LP3 that have similar frameworks to TRPO [29], TD($\lambda$) method [20] is used to train the distributional critic networks. For PD-MORL and CAPQL, the distributional critic networks are trained by reducing the Wasserstein distance between the current quantiles and the truncated one-step TD targets, as in TQC [21]. The policy network, similar to other RL algorithms dealing with continuous action spaces [29, 15, 14], outputs the mean and diagonal variance of a normal distribution. However, in the locomotion tasks, the policy networks are modified to output only the mean value by fixing the diagonal variance in order to lower the training difficulty.
C.3.3 Hyperparameter Settings
@@
$g_{\mathrm{min}}$, $g_{\mathrm{max}}$ $(0,\infty)$
$\lambda_{\mathrm{max}}$infty
$H$ matrix - - Fisher information Fisher information
@@
Policy learning rate $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
@@
Local region size - - 0.14 0.05
$g_{\mathrm{min}}$, $g_{\mathrm{max}}$ $(0,\infty)$
$\lambda_{\mathrm{max}}$infty
$H$ matrix - - Fisher information Identity matrix
Appendix D Additional Experimental Results
@@
In this section, estimated CP fronts for tasks with two objectives are visualized. As mentioned in the appendix C.2, the estimated CP front can be obtained by rolling out a trained universal policy with different preferences, and in this section, 20 equally spaced preferences are used. In Figure 12, the CP fronts for the legged robot locomotion tasks are visualized, Figure 13 presents the CP fronts for the Safety Gymnasium tasks, and Figure 14 presents the CP fronts for the MO-Gymnasium tasks. The hopper and ant tasks in the MO-Gymnasium are excluded from this visualization as they have three objectives.
(a) Biped
(b) Quadruped
Figure 12: Visualization of estimated CP fronts for legged robot locomotion tasks. For each task, we plot the estimated fronts obtained from five random seeds in the same figure.
(a) Single-Point-Goal
(b) Single-Car-Goal
@@
Figure 13: Visualization of estimated CP fronts for Safety Gymnasium tasks.
(a) Humanoid
(b) Half-Cheetah
(c) Walker2d
(d) Swimmer
Figure 14: Visualization of estimated CP fronts for MO Gymnasium tasks.